# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ Abstract
Achieving fully autonomous driving with heightened safety and efficiency depends on vehicle-to-everything (V2X) cooperative perception (CP), which allows vehicles to share perception data, thereby enhancing situational awareness and overcoming the limitations of the sensing ability of individual vehicles. V2X CP is crucial for extending perception range, improving accuracy, and strengthening the decision-making and control capabilities of autonomous vehicles in complex environments. This paper provides a comprehensive survey of recent advances in V2X CP, introducing mathematical models of CP processes across various collaboration strategies. We examine essential techniques for reliable perception sharing, including agent selection, data alignment, and fusion methods. Key issues are analyzed, such as agent and model heterogeneity, perception uncertainty, and the impact of V2X communication constraints like delays and data loss on CP effectiveness. To inspire further advancements in V2X CP, we outline promising avenues, including privacy-preserving artificial intelligence (AI), collaborative AI, and integrated sensing frameworks, as pathways to enhance CP capabilities.
@@
Autonomous driving; connected and automated vehicles (CAVs); cooperative perception (CP); V2X Communication; segmentation; object detection; multi-object tracking; data fusion; deep learning; artificial intelligence (AI).
@@ I Introduction
Perception plays a pivotal role in ensuring safety in autonomous driving, facilitating awareness of the surrounding road environment and other road participants over time, and utilizing various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8]. Recent advancements in intelligent vehicle technology have led to significant progress in single-vehicle perception techniques across domains, such as 3D object detection using camera [9, 10, 11, 12, 13], LiDAR [14, 15, 16, 17, 18, 19], radar [20, 21, 22, 23, 24, 25], and fusion [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], as well as 3D multi-object tracking employing camera [42, 43, 44, 45], LiDAR [46, 47, 48, 49, 50, 51], radar [52, 53, 54, 55], and fusion [56, 57, 58, 59, 60, 61, 62, 63], to name just a few. However, single-vehicle perception technologies still grapple with notable limitations. Sparse observations due to occlusion, restricted sensor Field of View (FoV), and sensor resolution remain primary challenges [64]. Additionally, susceptibility to sensor errors arising from adverse weather and hardware failures raises concerns about system robustness [65]. These bottlenecks constrain the potential and safety of automated vehicles in complex driving scenarios.
@@
Cooperative Perception (CP) has been developed to mitigate the constraints associated with the perception capabilities of individual vehicles. Enabled by Vehicle-to-Everything (V2X) communication [66], CP allows Connected Autonomous Vehicles (CAVs) to exchange information with each other (V2V) [67], infrastructure (V2I) [68, 69], pedestrians (V2P) [70], and networks (V2N) [71]. This communication facilitates the sharing of crucial information, such as traffic conditions, road hazards, and signal status, extending beyond the perceptual range of individual vehicles [72]. Through this data exchange, CP enhances decision-making, enabling vehicles to better anticipate the movements of other road users, synchronize actions, and execute coordinated maneuvers, such as platooning and safe lane changes.
In full-stack autonomous driving, components like perception, localization, planning, control, and system integration collectively enable a vehicle to operate autonomously and safely. V2X CP significantly enhances this stack by sharing sensory and positional data via V2X communication between connected vehicles and infrastructure, thereby extending each vehicle’s awareness beyond the limits of its sensing range. Through V2X CP, the perception layer benefits from improved accuracy and safety, as vehicles can detect occluded or distant objects using shared data from other agents111“Agents” refer to autonomous entities within traffic environments, including vehicles and infrastructure equipped with sensors and communication systems, working together to enable cooperative perception. in the environment. In localization, V2X CP refines vehicle positioning by leveraging positional data from surrounding agents, which is especially valuable in areas with poor GPS signals. Planning also becomes more adaptive and safer, as V2X CP allows access to real-time data on nearby vehicles’ movements, enabling vehicles to make smoother, more informed maneuvers. Additionally, the enhanced situational awareness that V2X CP provides indirectly improves control precision and supports seamless system integration, reducing the vehicle’s reliance on any single sensor. Overall, V2X CP strengthens the robustness and adaptability of autonomous systems, laying a foundation for safer and more scalable deployment in real-world scenarios.
Fig. 1 illustrates an urban scenario of CP supported by V2X communication. In Fig. 1a, each vehicle operates independently, relying solely on its perception systems without collaboration, which restricts their understanding of the environment with a limited sensing range. Due to obstructions and varying distances, autonomous vehicles face challenges such as unreliable perception of distant objects and obstructed views [73]. In contrast, Fig. 1b shows vehicles and the infrastructure collaborating through wireless communication, leading to a holistic understanding of the environment and eliminating blind spots. A specific scenario in Fig. 2 demonstrates a white hatchback whose front view is obstructed by a parked food truck, preventing it from detecting a boy chasing a football near the road, as depicted in Fig. 2a. This presents a significant collision risk if the vehicle fails to stop promptly. However, with CP enabled, the blue CAV shares information about the boy’s movement with the white hatchback, allowing it to anticipate and decelerate preemptively, as shown in Fig. 2b, thereby avoiding a fatal accident.
While CP enables safer driving, numerous additional factors must be considered to bring V2X CP into practical application, aligning with the foundational technologies underpinning V2X CP. For example, the associated data privacy [74, 75, 76] and ethical concerns [77], the associated development of smart cities [78], urban infrastructure [79, 80], the associated standardization, regulation, and policy from governments and regulatory bodies [81, 82, 83], the associated trust issue in human-system interaction [84],
@@
[91] 2023 -- The challenges of implementing CP within V2X systems and the practical applicability of different CP methodologies in V2X environments. $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$
@@
This survey 2024 -- The evolution of V2X CP technologies, key elements of V2X CP, and V2X CP solutions that tackle major issues in practical driving scenarios. $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
@@
and the balance between excessive reliance and information inundation [94]. Thus, while the potential of V2X CP is profound, a concerted, multi-dimensional effort is requisite for its holistic integration into our daily lives.
Although various approaches have been proposed to address these challenges, a comprehensive survey is needed to assess recent advancements, identify research gaps, and propose future directions. Some existing surveys often focus on specific aspects of CP, such as feature-level cooperative object detection [85] or infrastructure-based CP [86], but lack a systematic overview of V2X CP technologies. To fill this gap, the work in [87] provided an overview of CP algorithms, emphasizing challenges like localization and detection, while the review in [88] discussed communication delays and information fusion in multi-sensor V2V and V2I schemes. The survey in [89] explored CP system architectures and proposed a hierarchical framework for unifying CP scenarios. Despite covering common CP challenges, these surveys lacked an analysis of CP progress regarding practical implementation challenges.
Recent work [90] evaluated collaborative modules and solutions focusing on efficiency, robustness, and safety. Another survey [91] addressed practical challenges in implementing CP, such as latency, bandwidth constraints, and robustness, categorizing methodologies into early, intermediate, and late fusion. Environmental simulations were also conducted to assess mainstream CP approaches. A more recent survey [92] focused on CP at intersections, analyzing perception, communication, and downstream applications. However, these surveys overlook the critical role of V2X communication, which is essential for enabling effective CP in autonomous vehicles by facilitating information transfer and supporting decision-making. The work in [93] examined typical V2X communication technologies, including Dedicated Short-range Communication (DSRC) and Cellular-based V2X communication (C-V2X), and presented an overview of perception architectures and technologies within the Vehicle-Road-Cloud Integration System, encompassing single-node, multi-node, and vehicle-road-cloud cooperative perception frameworks. While it offers insights into these technologies, it does not explore specific algorithms and techniques in depth. This work contributes to a comprehensive understanding of CP and its enabling V2X technologies and builds upon existing survey efforts. Table I compares our paper with these surveys. The key contributions are as follows:
TABLE II: List of Key Acronyms.
Acronyms Definitions
AI Artificial Intelligence
BEV Bird’s Eye View
CAM Cooperative Awareness Message
CAV Connected Autonomous Vehicle
CDA Cooperative Driving Automation
CI Covariance Intersection
CNN Convolutional Neural Network
CP Cooperative or Collaborative Perception
DSRC Dedicated Short-range Communication
C-V2X Cellular-based Vehicle-to-Everything
IRS Intelligent Reflecting Surface
CPM Collective Perception Message
DT Digital Twin
FL Federated Learning
FoV Field of View
GNN Graph Neural Network
GNSS Global Navigation Satellite System
GPS Global Positioning System
IMU Inertial Measurement Unit
ISAC Integrated Sensing and Communication
ITS Intelligent Transportation System
LSTM Long Short-term Memory
MIMO Multiple-input Multiple-output
RSU Roadside Unit
V2I Vehicle-to-Infrastructure
V2V Vehicle-to-Vehicle
V2X Vehicle-to-Everything
  • Refer to caption
    Figure 3: Organization of this paper.
@@
The structure of this paper is illustrated in Fig. 3. Section II presents the mathematical formulation of V2X CP. A brief historical overview of CP system development is provided in Section III. Section IV introduces a generic framework for systematically understanding the CP system. Practical issues related to CP systems and relevant datasets are discussed in Sections V and VI, respectively. Section VII compares the performance of various CP methods in offline testing and explores real-world implementations. Section VIII explores the challenges and potential future directions, and the paper concludes with Section IX. Table II lists key acronyms and abbreviations used throughout the paper.
@@ II Mathematical Formulation of Cooperative Perception
This section begins with an overview of sensors and CP collaboration strategies, followed by a detailed mathematical formulation for early, intermediate, and late collaboration strategies.
II-A Overview
CP is built upon the shared data contributed by each participating agent. For each individual agent, raw observations are collected via onboard sensors, including cameras, LiDAR, radar, and, in some cases, ultrasonic sensors. Cameras provide rich visual information—such as color, texture, and shape—essential for object detection and scene interpretation [95]. LiDAR generates high-resolution 3D point clouds by measuring the return time of laser pulses, enabling precise distance and shape estimation critical for detailed mapping and obstacle detection [96]. Radar supplies range, speed, and movement data, effectively detecting objects at greater distances and under diverse weather conditions, making it especially valuable for tracking the velocity of surrounding vehicles. Although ultrasonic sensors are less commonly utilized in CP, they contribute valuable proximity data in close-range applications, such as parking and low-speed maneuvers.
CP strategies are categorized into three collaboration types: early, late, and intermediate collaboration, each defined by the level of data processing before sharing. Note that all types of sensor data can be utilized across any of these collaboration types. In early collaboration, raw sensor data is shared directly among agents, providing high-fidelity information but requiring substantial communication bandwidth and precise synchronization. This approach can be challenging in networks with limited bandwidth or high latency, leading to potential inconsistencies in data across agents. In late collaboration, only the final perception outputs, such as object classifications and location, are shared. This minimizes communication demands significantly but limits data granularity, potentially restricting the system’s ability to perform detailed analysis or adapt to complex, rapidly evolving scenarios. Intermediate collaboration strikes a balance by pre-processing sensor data to extract features before sharing them with other agents. This approach reduces data volume while retaining essential information, lowering bandwidth requirements. However, the effectiveness depends on each agent’s feature extraction accuracy, and variability in feature quality may lead to perception inconsistencies across agents.
A schematic overview of these collaboration schemes is provided in Fig. 4. For mathematical clarity, a scenario with $N$ collaborating agents is considered, where the local raw sensor data observed by the $i$-th agent at time step $t$ is denoted as $\boldsymbol{X}_{i}^{t}$. In the following subsections, the workflow and mathematical formulation of each collaboration scheme are elaborated in detail.
@@ II-A Overview
(b) Late
(c) Intermediate
Figure 4: Overview of collaboration schemes for autonomous driving: (a) Early collaboration; (b) Late collaboration; (c) Intermediate collaboration.
II-B Early Collaboration
As illustrated in Fig. 4a, early collaboration involves aggregating raw sensor observations collected by all participating agents, fostering a comprehensive perspective [97, 98, 99]. The $i$-th agent broadcasts its raw observation $\boldsymbol{X}_{i}^{t}$ to the other $N-1$ neighbouring agents and also receives the raw observation $\boldsymbol{X}_{j}^{t},j\in\{1,\cdots,N\},j\neq i$. After that, it performs data fusion on the received raw data with self-raw data and produces fused CP raw data $\boldsymbol{O}_{i}^{t}$. For simplicity, we assume that all agents use the same fusion, encoder, and decoder algorithms. Thus, the raw data fusion at the $i$-th agent can be expressed as:
$\boldsymbol{O}_{i}^{t}=\mathcal{F}_{\mathrm{EarlyFusion}}\left({\left\{\boldsymbol{X}_{k}^{t}\right\}}_{k=1}^{N}\right),$ (1)
where $\mathcal{F}_{\mathrm{EarlyFusion}}(\cdot)$ denotes the early fusion algorithm.
Then, the $i$-th agent uses an encoder $\mathcal{F}_{\mathrm{Encoder}}(\cdot)$ to extract high-level CP feature $\boldsymbol{\widetilde{F}}_{i}^{t}$ from its fused CP raw data $\boldsymbol{O}_{i}^{t}$:
$\boldsymbol{\widetilde{F}}_{i}^{t}=\mathcal{F}_{\mathrm{Encoder}}(\boldsymbol{O}_{i}^{t}).$ (2)
Finally, the $i$-th agent uses a decoder $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ to decode the CP feature $\boldsymbol{\widetilde{F}}_{i}^{t}$ and obtain the CP results $\widetilde{\boldsymbol{Y}}_{i}^{t}$:
$\widetilde{\boldsymbol{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{\widetilde{F}}_{i}^{t}\right).$ (3)
II-C Late Collaboration
In contrast to early collaboration, late collaboration involves exchanging individual perceptual outcomes, as depicted in Fig. 4b, requiring minimal data transmission [100].
In this case, the $i$-th agent independently processes its raw perception data $\boldsymbol{X}_{i}^{t}$, generate its feature data $\boldsymbol{F}_{i}^{t}$ and its final perception result $\boldsymbol{R}_{i}^{t}$, via (4) and (5) respectively:
$\boldsymbol{F}_{i}^{t}=\mathcal{F}_{\mathrm{Encoder}}\left(\boldsymbol{X}_{i}^{t}\right),$ (4)
$\boldsymbol{R}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{{F}}_{i}^{t}\right),$ (5)
where $\mathcal{F}_{\mathrm{Encoder}}(\cdot)$ and $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ represent the encoder and decoder algorithms, respectively.
After individual processing, the individual perception results from each agent are shared with all other agents. Then, for the $i$-th agent, it uses a late fusion algorithm $\mathcal{F}_{\mathrm{LateFusion}}(\cdot)$ to fuse the received individual perception result from all other $N-1$ agents and its own perception result to generate CP result $\boldsymbol{\widehat{Y}}_{i}^{t}$:
$\boldsymbol{\widehat{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{LateFusion}}\left({\left\{\boldsymbol{R}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (6)
II-D Intermediate Collaboration
In this case, as shown in Fig. 4c, each agent shared intermediate features rather than raw data or final perception results. Firstly, the $i$-th agent extracts its own feature data $\boldsymbol{F}_{i}^{t}$ from its raw perception data $\boldsymbol{X}_{i}^{t}$ as described in (4). Then, the extracted individual feature data are shared among all $N$ agents via V2X communication. After receiving shared feature data from other $N-1$ agents, the $i$-th agent use an intermediate fusion algorithm $\mathcal{F}_{\mathrm{InterFusion}}(\cdot)$ to generate the fused CP feature data $\boldsymbol{H}_{i}^{t}$:
$\boldsymbol{H}_{i}^{t}=\mathcal{F}_{\mathrm{InterFusion}}\left({\left\{\boldsymbol{F}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (7)
Finally, the fused CP feature $\boldsymbol{H}_{i}^{t}$ is fed into the decoder $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ to obtain the final CP result ${\boldsymbol{Y}}_{i}^{t}$ at the $i$-th agent:
${\boldsymbol{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{H}_{i}^{t}\right).$ (8)

### v5

# Vehicle-to-Everything Cooperative Perception for Autonomous Driving

@@ Abstract
Achieving fully autonomous driving with enhanced safety and efficiency relies on vehicle-to-everything cooperative perception, which enables vehicles to share perception data, thereby enhancing situational awareness and overcoming the limitations of the sensing ability of individual vehicles. Vehicle-to-everything cooperative perception plays a crucial role in extending the perception range, increasing detection accuracy, and supporting more robust decision-making and control in complex environments. This paper provides a comprehensive survey of recent developments in vehicle-to-everything cooperative perception, introducing mathematical models that characterize the perception process under different collaboration strategies. Key techniques for enabling reliable perception sharing, such as agent selection, data alignment, and feature fusion, are examined in detail. In addition, major challenges are discussed, including differences in agents and models, uncertainty in perception outputs, and the impact of communication constraints such as transmission delay and data loss. The paper concludes by outlining promising research directions, including privacy-preserving artificial intelligence methods, collaborative intelligence, and integrated sensing frameworks to support future advancements in vehicle-to-everything cooperative perception.
@@
Autonomous driving; connected and automated vehicles; cooperative perception; vehicle-to-everything communication; segmentation; object detection; multi-object tracking; data fusion; deep learning; artificial intelligence.
@@ I Introduction
Cooperative Perception (CP), enabled by Vehicle-to-Everything (V2X) communication [1], has emerged as a promising paradigm for enhancing situational awareness in autonomous driving. By enabling vehicles to exchange perception data with surrounding vehicles, infrastructure, pedestrians, and networks, CP extends visibility beyond line-of-sight limitations and enhances decision-making in complex scenarios, such as intersections and lane changes. This collaborative approach addresses the fundamental limitations of single-vehicle perception, which depends solely on onboard sensors such as cameras, LiDAR, and radar.
In a full-stack autonomous driving system, core components, including perception, localization, planning, control, and system integration, must operate cohesively. V2X-based CP enhances each of these layers by enabling the exchange of sensory and positional data among connected agents111“Agents” refer to vehicles and infrastructure equipped with sensing and communication capabilities.. At the perception level, shared information allows the detection of objects beyond a vehicle’s sensing range, improving overall accuracy. In localization, data from neighboring agents complements Global Positioning System (GPS) signals, particularly in areas with poor satellite coverage. For planning, real-time awareness of nearby trajectories enables more adaptive and informed decisions. These improvements also benefit the control layer, where enhanced situational awareness leads to more precise execution. Finally, system integration is strengthened through reduced reliance on individual sensors, enhancing the platform’s robustness and scalability. Collectively, V2X CP improves the intelligence and resilience of autonomous systems, supporting their safe deployment in real-world scenarios.
@@
Fig. 1 illustrates an urban scenario comparing autonomous driving with and without CP. In Fig. 1a, each vehicle operates independently. Each vehicle’s limited sensing range, combined with obstructions, leads to unreliable detection of distant or occluded objects [2]. In contrast, Fig. 1b depicts vehicles and infrastructure exchanging information via V2X communication, resulting in a more comprehensive understanding of the environment and the elimination of blind spots. A specific case is shown in Fig. 2, where a white hatchback’s front view is blocked by a parked food truck, preventing it from detecting a child running into the street. This scenario, also shown in Fig. 2a, poses a serious collision risk if the vehicle cannot respond in time. With CP enabled, the blue Connected Autonomous Vehicle (CAV) communicates the child’s movement to the white hatchback, allowing it to decelerate proactively and avoid an accident, as illustrated in Fig. 2b.
While CP enhances driving safety, its practical deployment requires addressing several foundational challenges, including data privacy [3], ethical concerns [4], integration with smart cities and urban infrastructure [5], and regulatory and policy frameworks [6]. Realizing the full potential of V2X CP, therefore, demands a coordinated, multi-dimensional effort.
@@
[13] 2023 - The challenges of implementing CP within V2X systems and the practical applicability of different CP methodologies in V2X environments. $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$
@@
This survey 2025 - The evolution of V2X CP technologies, key elements of V2X CP, and V2X CP solutions that tackle major issues in practical driving scenarios. $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
@@
Although various approaches have been proposed to address these challenges, a comprehensive survey is needed to assess recent advances, identify research gaps, and propose future directions. Some existing surveys often focus on specific aspects of CP, such as feature-level cooperative object detection [7] or infrastructure-based CP [8], but lack a systematic overview of V2X CP technologies. To fill this gap, the work in [9] provided an overview of CP algorithms, emphasizing challenges like localization and detection, while the review in [10] discussed communication delays and information fusion in multi-sensor Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) schemes. The survey in [11] explored CP system architectures and proposed a hierarchical framework for unifying CP scenarios. Despite covering common CP challenges, these surveys lacked an analysis of CP progress regarding practical implementation challenges.
Recent work [12] evaluated collaborative modules and solutions focusing on efficiency, robustness, and safety. Another survey [13] addressed practical challenges in implementing CP, such as latency, bandwidth constraints, and robustness, categorizing methodologies into early, intermediate, and late fusion. Environmental simulations were also conducted to assess mainstream CP approaches. A more recent survey [14] focused on CP at intersections, analyzing perception, communication, and downstream applications. However, these surveys overlook the critical role of V2X communication, which is essential for enabling effective CP in autonomous vehicles by facilitating information transfer and supporting decision-making. The work in [15] examined typical V2X communication technologies, including Dedicated Short-range Communication (DSRC) and Cellular-based V2X communication (C-V2X), and presented an overview of perception architectures and technologies within the Vehicle-Road-Cloud Integration System, encompassing single-node, multi-node, and vehicle-road-cloud cooperative perception frameworks. While it offers insights into these technologies, it does not explore specific algorithms and techniques in depth. This work contributes to a comprehensive understanding of CP and its enabling V2X technologies and builds upon existing survey efforts. Table I compares our paper with these surveys. The key contributions of this paper are as follows:
@@
The remainder of this paper is organized as follows: Section II presents the mathematical formulation of V2X CP. A brief historical overview of CP system development is provided in Section III. Section IV introduces a generic framework for systematically understanding the CP system. Practical issues related to CP systems and relevant datasets are discussed in Sections V and VI, respectively. Section VII compares the performance of various CP methods in offline testing and explores real-world implementations. Section VIII explores the challenges and potential future directions, and the paper concludes with Section IX.
@@ II Mathematical Formulation of Cooperative Perception
CP strategies can be categorized into three types based on the processing stage at which data is shared: early, intermediate, and late collaboration. Let $\boldsymbol{X}_{i}^{t}$ denote the raw sensor data from agent $i$ at timestep $t$. Each agent $i$ uses an encoder, $\mathcal{F}_{\mathrm{Encoder}}(\cdot)$, to extract features $\boldsymbol{F}_{i}^{t}$ from raw sensor data, and then a decoder, $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$, to generate perception results $\boldsymbol{Y}_{i}^{t}$. The key distinction across strategies lies in the data sharing stage and corresponding fusion method. A schematic overview of these collaboration schemes is provided in Fig. 3.
II-A Early Collaboration
In early collaboration, agents exchange raw sensor data. Each agent $i$ fuses data from all $N$ agents:
$\boldsymbol{X}_{i,\mathrm{CP}}^{t}=\mathcal{F}_{\mathrm{EarlyFusion}}\left({\left\{\boldsymbol{X}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (1)
The fused raw data $\boldsymbol{X}_{i,\mathrm{CP}}^{t}$ is then encoded and decoded at agent $i$ to produce the final CP output. While this method provides the richest information, it imposes high bandwidth and synchronization requirements, making it less practical in communication-constrained environments.
II-B Intermediate Collaboration
Intermediate collaboration involves sharing features extracted from raw data. Each agent $i$ encodes its observation as a feature $\boldsymbol{F}_{i}^{t}$ and exchanges it with peers. Fusion is performed on the feature level:
$\boldsymbol{F}_{i,\mathrm{CP}}^{t}=\mathcal{F}_{\mathrm{InterFusion}}\left({\left\{\boldsymbol{F}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (2)
The fused feature information $\boldsymbol{F}_{i,\mathrm{CP}}^{t}$ is then decoded at agent $i$ to obtain the CP output. This strategy balances communication efficiency and information richness, but its performance depends on the quality and consistency of features across agents.
II-C Late Collaboration
In late collaboration, each agent $i$ completes local perception independently and shares only the final output $\boldsymbol{Y}_{i}^{t}$ with other agents. These results are aggregated using a late fusion function:
$\boldsymbol{Y}_{i,\mathrm{CP}}^{t}=\mathcal{F}_{\mathrm{LateFusion}}\left({\left\{\boldsymbol{Y}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (3)
This approach minimizes bandwidth requirements but lacks flexibility and detail, making it less suitable for dynamic or occluded environments.
@@ II-C Late Collaboration
(b) Intermediate
(c) Late
Figure 3: Overview of collaboration schemes for autonomous driving: (a) Early collaboration; (b) Intermediate collaboration; (c) Late collaboration.